# Time-dependent Hamiltonians and Geometry of Operators Generated by Them

### v1

@@ Appendix A Quantum quenches - an exact example of time-dependent scenarios
A.1 Change in the Shannon entropy
@@
$\rho(t)=\sqrt{\frac{\eta}{\beta}\bigg{(}\frac{s^{2}+1}{1+\beta^{2}}\bigg{)}}\bigg{(}\cos^{2}\Big{[}\sqrt{1+\beta^{2}}\arctan[s]\Big{]}+\beta^{2}\bigg{)}^{1/2}~{}.$ (113)
@@
A.2 Cost of the time evolution operator
Next, we compute the cost of the time evolution operator for this frequency profile with the help of the formula derived in section IV and study its various limits. The expression for the time-dependent function $f(t)$, obtained by using the solution of the auxiliary function written above (see Eq. (113)) into Eq. (40), is given in terms of the dimensionless time $s$  by the formula
@@ A.2 Cost of the time evolution operator
Figure 3: Variation of the cost ($F_{N}^{2}(s)=\eta^{2}F_{r}^{2}(s)$) with respect to $\beta$ for different values of $s$. Here the fixed values of the dimensional time are $s=2$ (red), $s=20$ (green), and $s=200$ (orange).
Figure 4: Scaling of the cost ($F_{N}^{2}(s)=\eta^{2}F_{r}^{2}(s)$) at late times for fast quenches. The dashed green curve represents the approximate cost obtained using Eq. (116), while the solid red is the exact one. Quadratic and linear scaling are also shown by dot-dashed black and brown fittings. Here, we have fixed dimensional time $s=200$.
We can obtain a scaling behaviour of the cost at late times ($s\gg 1$) and for $\beta\ll 1$ (i.e., for the case-2 considered for the SE above). Using the expansion of the auxiliary function given in Eq. (116) for late times and small $\beta$, in the expression for the cost, we can obtain an approximate variation of $F_{N}^{2}(s)$ in this limit. This is plotted in Fig. 4 (the dashed green curve), along with the exact expression (red plot), and as can be seen, the truncated expression provides a good approximation for the cost function. Furthermore, in this plot, we have also shown the initial quadratic scaling of the cost for small $\beta$, which later matches with a linear scaling. For the fittings shown in Fig. 4, the numerically obtained equations for the curves are: quadratic scaling $F_{N}^{2}=\beta^{2}/18$ (shown by the dot-dashed black curve), and for linear scaling $F_{N}^{2}=0.0345(\beta-0.17)$ (shown by dot-dashed brown line).
@@
Appendix B Decomposition formula for the time evolution operator of an OTMF

### v2

@@ VI Quantum quenches - an exact example of time-dependent protocols
VI.1 Change in the Shannon entropy
@@
$\rho(t)=\sqrt{\frac{\eta}{\beta}\bigg{(}\frac{s^{2}+1}{1+\beta^{2}}\bigg{)}}\bigg{(}\cos^{2}\Big{[}\sqrt{1+\beta^{2}}\arctan(s)\Big{]}+\beta^{2}\bigg{)}^{1/2}~{}.$ (101)
@@
VI.2 Cost of the time evolution operator
Next, we compute the cost of the time evolution operator for this frequency profile with the help of the formula derived in section IV and study its various limits. The expression for the time-dependent function $f(t)$, obtained by using the solution of the auxiliary function written above (see Eq. (101)) into Eq. (30), is given in terms of the dimensionless time $s$  by
@@ VI.2 Cost of the time evolution operator
Figure 3: Variation of the cost ($F_{N}^{2}(s)=\eta^{2}F_{r}^{2}(s)$) with respect to $\beta$ for different values of $s$. Here the fixed values of the dimensional time are $s=2$ (red), $s=20$ (green), and $s=200$ (orange). For small values of $\beta$ ($10<\beta<$ 1) the cost always grows irrespective of the value $s$ and subsequently reaches a local peak.
Figure 4: Scaling of the cost ($F_{N}^{2}(s)=\eta^{2}F_{r}^{2}(s)$) with respect to $\beta$ at late times for fast quenches. The dashed green curve represents the approximate cost obtained using Eq. (104), while the solid red is the exact one. Quadratic and linear scaling are also shown by dot-dashed black and brown fittings. Here, we have fixed dimensional time $s=200$.
We can obtain a scaling behaviour of the cost at late times ($s\gg 1$) and for $\beta\ll 1$ (i.e., for the case-2 considered for the SE above). Using the expansion of the auxiliary function given in Eq. (104) for late times and small $\beta$, in the expression for the cost, we can obtain an approximate behaviour of $F_{N}^{2}(s)$ in this limit. This is plotted in Fig. 4 (the dashed green curve), along with the exact expression (red), and as can be seen, the truncated expression provides quite a good approximation for the cost function. Furthermore, in this plot, we have also shown the initial quadratic scaling of the cost for small $\beta$, which subsequently transforms with a linear scaling. For the fittings shown in Fig. 4, the numerically obtained equations for the scalings are: quadratic scaling $F_{N}^{2}=\beta^{2}/18$ (shown by the dot-dashed black curve), and for linear scaling $F_{N}^{2}=0.0345(\beta-0.17)$ (shown by dot-dashed brown line).
@@
VII Summary and discussions
In this paper we have discussed a method of obtaining the regularised bi-invariant CG of the operator generated by the time-dependent Hamiltonian of a quantum mechanical system. The procedure we have followed makes use of a regularisation scheme in which the norm of an infinite dimensional Hamiltonian is obtained by first imposing a condition on its trace and subsequently using the resultant mean value ($a$) in the calculation of the cost function. For time-independent Hamiltonians (which we had considered in section III), this mean value is also a constant, and therefore, the cost itself is independent of time. However, for explicitly time-dependent Hamiltonians considered in later sections, this mean value is itself a function of time, and as a result, the computation of the regularised cost is more involved. Another feature of the calculation where the computation of the cost of the two types of systems differs is that, for time-independent Hamiltonians, one only needs to know the energy eigenvalues to calculate the cost, whereas since a time-dependent Hamiltonian has no stationary eigenstates, both diagonal as well as non-diagonal matrix elements of this Hamiltonian in a suitable basis are required for the computation of the cost. To this end, we have used the well known Lewis-Riesenfeld method of time-dependent invariant operators since the eigenfunctions of this operator can be directly used efficiently to find out both the diagonal and non-diagonal matrix elements of the Hamiltonian under consideration, and hence the norm can be calculated in terms of these matrix elements. This is the strategy we have followed in section IV and V to obtain the regularised norm of time-dependent Hamiltonians, where the generic properties of the obtained results are also discussed.
As emphasised in the main text, we again want to stress the importance of looking at the standard problem of circuit complexity from a slightly different viewpoint: rather than finding a minimum of all unitary curves connecting the two given operators (which is the standard Nielsen’s geometric framework of obtained the circuit complexity), calculating the total cost of evolution along a specified path on a CG is also important, since the time-dependent Hamiltonian of a quantum mechanical system may itself generate such unitary evolution. In that situation, it is meaningful only to compute the cost as a function of time along that specified evolution curve, not the complexity. The assumption of bi-invariance of the CG, and the regularisation procedure used to find out the norm for an infinite dimensional Hilbert space, are two key ingredients of our analysis, and both have strong effects on the final form of the cost function we have derived for different classes of harmonic oscillators.
Among various examples of explicitly time-dependent systems, in this paper, we have focused on two particular Hamiltonians, which are widely studied in the literature due to their importance in modelling various physical phenomena. These are, respectively, an oscillator of unit mass whose frequency of oscillation explicitly depends on time and an oscillator whose both mass and frequency change with time. For convenience here we briefly list the main results we have obtained in this paper along with some of their implications: 1. For an OTF and an OTMF, using the Lewis-Riesenfeld method, we have obtained compact analytical expressions for the cost of the time evolution operator generated by these two Hamiltonians (see eqs. (47) and (74), respectively). 2. We have shown there exists an equivalence between the two separate examples considered so that expression for various quantities of two systems can be mapped to one another by a time reparametrisation, and demonstrated the implications of this duality in terms of the total bi-invariant cost. 3. We have obtained an upper bound on the total cost up to a finite time (see eqs. (51) and (76)), which is analogous to the quantum speed limits associated with various quantum information-theoretical quantities (usually a quantifier of ‘distance’ between two quantum states). 4. One of the most important examples of time-dependent quantum systems, namely a smooth quench of the Hamiltonian parameters is considered in section VI. For a time-dependent frequency oscillator model, using the exact analytical expressions available in the literature, we have studied information theoretic quantities, e.g., change of the Shannon entropy, and compared them with the cost of the time-evolution operator derived here. In both cases, we have shown that different quench regimes and their associated characteristics can be effectively probed by these quantifiers by using their scaling relations.
We also now briefly discuss the differences between the cost function we have derived in this paper, and the approach followed in general for the case of a time-dependent Hamiltonian, with the existing results in the literature. The approach developed in Myers1 ; BSS to compute the NC in quantum field theories, which is one of the most common approaches followed by many authors, directly uses the wavefunction of the state which differ from each other in terms of the values of the parameter of the system (e.g., the mass and the frequency for a harmonic oscillator), not time, and is therefore, not directly related to the time-evolution of the state. On the other hand, refs. BDKP1 ; BDKP2 studied NC of time-evolution operator. However, the role played by the time parameter in BDKP1 ; BDKP2 is quite different from the one in this paper, since, in the approach followed by these authors, the parameter that connects the initial and final points on the CG is different from the time. Therefore, in our opinion, it is not very meaningful to directly compare our results for the cost function with the complexity that appeared so far in the literature, simply because, in general, they refer to different quantities, albeit in the same geometrical framework.
Before concluding, we mention a few possible future directions. Firstly, in this paper we have mainly considered the generalised oscillator and other Hamiltonians which can be connected to it by unitary transformations. Extending our analysis to systems beyond this class, specifically systems that are classically chaotic, can be one important direction that can be pursued. Secondly, the bi-invariant CG we have discussed in this paper is invariant under unitary transformations between the generators which are independent of the parameter describing the unitary curve, e.g., as we have discussed in section V.1, the unitary transformation $U_{1}$ in (62) connecting the Hamiltonians of a generalised oscillator and an OTMF is time-independent, hence, the cost is the same for both the Hamiltonians. However, if we consider two Hamiltonians which are related through a time-dependent unitary transformation, the CG of Eq. (11) is no longer invariant under such transformation (this is the reason why the costs associated with OTF and OTMF are not the same). Determining an invariant CG generated by Hamiltonians related through general time-dependent unitary transformations is an interesting problem, and we hope to come back to it in a future work.
Acknowledgments
We sincerely thank Hugo A. Camargo, Yichao Fu, Viktor Jahnke, Keun-Young Kim, and Run-Qiu Yang for useful discussions on this and other related works, as well as comments on a draft version of the manuscript. We would also like to thank the organisers of the workshop “Student Talks in Trending Topics in Theory” (ST4 -2023), during which parts of the present work were carried out. The work of Kunal Pal is supported by the National Research Foundation of Korea under Grants No. 2017R1A2A2A05001422 and No. 2020R1A2C2008103. We thank our anonymous referees for their constructive comments and criticisms which helped to improve a draft version of this manuscript.
Appendix A Further examples of bi-invariant complexity geometry with time-independent Hamiltonian
A.1 Complexity geometry generated by the generalised oscillator Hamiltonian with a constant driving force
which, as the subscript above suggests, is just that of a HO with a constant term added (here $B_{1}=B_{0}$). Notice that this constant term actually shifts the energy eigenvalues of the system, so that the analogues of the relations in Eq. (12) and Eq. (22) can be written without introducing the constant $a$ by hand, and are given by,
A.2 Complexity geometry generated by the isotonic oscillator Hamiltonian
The norm of the Hamiltonian, in turn, is given by
Appendix B Complexity geometry associated with the Caldirola–Kanai Hamiltonian
Appendix C Decomposition formula for the time evolution operator of an OTMF